ilings Review View Help Tell me what you want to do AaBbCcI AaBbCcl naBbC AaB Normal TNo Spac. 1 Headin Paragraph Styles Dally Cleaners reported the following information at December 31,2018: Common stock, $2 par, 300,000 shares authorized Additional paid-in capital Retained earnings 80,000 1,200,000 Total $6,880,000 68,000 Less: Treasury stock (2,000 common shares at cost) Total Stockholders Equity Net Income $845,000 Answer the following questions for Daly Cleaners in tke space provided 1. How many shares of common stock are authorized? 2. How many shares of common stock are issued 3. How many shares of common stock are outstanding? 4. How masy of the common shares will receive dnidends when they are paid? S. How much did Dally Cleaners pay for each saare of Treasury stock? 6. If a total of $44,300 diriderds ate paid, how much will shareholders receite per share What is tde total amount of coatributed capital at December 3, 201s? 7. . Calculas tae EPS (Tanings pe shxe for fo
Answer #1

1.Authorised shares = 300,000

2. Issued Shares = 80,000/2 = 40,000 shares

3.Outstanding Shares = Issued Shares – Treasury Stock

= 40,000 – 2,000 = 38,000 shares

4. Outstanding shares would receive dividend i.e. 38,000 shares

5.Price paid for each share of treasury stock = 68,000/2,000 = $34 per share

6.Dividend per share = $44,800/38,000 = $1.179 (approx.)

8. Earnings per share = 845,000/38,000 = $22.24 (approx.)

9.Price Earning Ratio = Price per Share/Earnings per Share